Adults consume less than about 25g of sugar a day, What percentage of this mass of sugar is found in 15g of drink crystals


1
Expert's answer
2021-10-05T23:56:23-0400

Drink crystals contain approximately 67 g of sugar per 100 g.

15 g of drink crystals contain 67/100 x 15 = 10.05 g of sugar.

This constitutes the folowing percentage from the daily sugar dose:

% = 10.05/25 x 100 = 40.2%